OpenStudy (anonymous):

Can i get help finding the derivative y = ln ((1+e^x)/(1-e^x))

OpenStudy (anonymous):

Use the chain rule, first find the derivative of 1+e^x/1-e^x, then multiply it by the derivative of ln z where z would be 1+e^x/1-e^x.

OpenStudy (anonymous):

could you possibly show me how to do the problem by expanding it first?

OpenStudy (anonymous):

ln(a/b)=ln a - ln b. So in this case: y=ln(1+e^x)-ln(1-e^x), then you use the chain rule again separetely on each term

OpenStudy (anonymous):

It gets: y'=e^x/(1+e^x)-(-e^x)/(1-e^x)=e^x/(1+e^x)+e^x/(1-e^x)=2e^x/(1-e^(2x))

OpenStudy (anonymous):

okay thats what i ended up getting thank you for making it much more clear for me